Determine the current flowing through an element if the charge flow is given by a. q(t)= (3e^-t -5e-^2t) nC b. q(t)= 10sin120πt pC c. q(t)= 20e^-4t cos50t µC

1
Expert's answer
2021-11-17T10:42:26-0500

(a) The current can be found as


"I=\text dq/\text d t,\\\space\\
I=\frac{\text d}{\text d t}(3e^{-t} -5e^{-2t})=e^{-2t}(10-3e^t)\text{ nA}."

(b) The current is


"I=\text dq/\text d t,\\\space\\
I=\frac{\text d}{\text d t}[10\sin(120πt)]=1200\pi\cos(120\pi t)\text{ pA}."

(c) The current:


"I=\frac{\text d}{\text d t}\Big[20e^{-4t}·\cos(50t)\Big]=\\\space\\
=-40e^{-4t}[25\sin(50t)+2\cos(50 t)]\space\mu\text{A}."